40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 3 18:49:44 2013, Gautam, Update, General, ITMx Oplev Fixed  
    Reply  Wed Jun 5 20:45:19 2013, gautam, Update, General, ITMx Oplev Fixed  ITMx_overview.JPGITMx_clipping_.JPG
       Reply  Wed Jun 26 19:02:38 2013, gautam, Update, General, ITMx Oplev (not quite?) Fixed oplev_chamber.pdfphoto_1.JPGphoto_2.JPG
          Reply  Thu Jun 27 18:11:53 2013, gautam, Update, General, ITMx Oplev-POX looks beam okay 
             Reply  Thu Jun 27 23:01:39 2013, gautam, Update, General, ITMx Oplev-servo gains adjusted pitch-plot.pdfyaw-plot.pdf
                Reply  Fri Jun 28 13:10:28 2013, gautam, Update, General, ITMx Oplev-servo gains adjusted pitch-plot_copy.pdfyaw-plot_copy.pdf
Message ID: 8684     Entry time: Wed Jun 5 20:45:19 2013     In reply to: 8665     Reply to this: 8758
Author: gautam 
Type: Update 
Category: General 
Subject: ITMx Oplev Fixed  

Quote:

 (Manasa)

The ITMx Oplev was misaligned. Switched the ITMx Oplev back on and fixed the alignment.

EDIT, JCD:  This is totally my fault, sorry.  I turned it off the other day when I was working on the POP layout, and forgot to turn the laser back on.  Also, I moved the fork on the lens directly in front of the laser (in order to accommodate one of the G&H mirrors), and I nudged that lens a bit, in both X and Y directions (although very minimally along the beam path).  Anyhow, bad Jenne for forgetting to elog this part of my work.

 [Jenne, Gautam]

It turned out that the earlier fix was not really a fix, because there was some confusion as to which of the two lenses Jenne moved while working, and while Manasa and I were re-aligning the beam, we may have moved the other lens.

Subsequently, when we checked the quadrant sum, it was low (in the region of 20), even though OPLEV_PERROR and OPLEV_YERROR were reasonably low. We called up a 30 day trend of the quadrant sum and found that it was typically closer to 4000. This warranted a visit to the table once again. Before going to the table, we did a preliminary check from the control room so as to make sure that the beam on the QPD was indeed the right one by exciting ITMx in pitch (we tried offsets of 500 and -500 counts, and the spot responded as it should). ITMx oplev servo was then switched off.

At the table, we traced the beam path from the laser and found, first, that the iris (I have marked it in one of the photos attached) was practically shut. Having rectified this, we found that the beam was getting clipped on the first steering mirror after the laser (also marked in the same photo, and a second photo showing the clipping is attached). The beam isn't very well centred on the first lens after the laser, which was the one disturbed in the first place. Nevertheless, the path of the entering beam seems alright. The proposed fix, then, is as follows;

  1. Use the existing iris and a second one to mark the path of the entering beam.
  2. Make the necessary adjustments (i.e. move the lens immediately after the laser and the first steering mirror after the laser) such that the direction of the entering beam is unchanged, and the clipping and centering problems are resolved. 

Back in the control room we noticed that the quadrant sum had gone up to ~3500 after opening out the iris. The OPLEV_PERROR and OPLEV_YERROR counts however were rather high (~200 counts in pitch and ~100 counts in yaw). Jenne went back to the table and fixed the alignment such that these counts were sub-10, and the quadrant sum went up to ~3800, close to the trend value. 

At the time of writing, the beam is still not centred on the lens immediately after the laser and is still getting clipped at the first steering mirror. Oplev servo back on.

Photos

 An overview

ITMx_overview.JPG

 

The clipping

ITMx_clipping_.JPG

ELOG V3.1.3-